Derek Jones, a foreign exchange trader at Charles Schwab, can invest $1 million, or the foreign currency equivalent of the bank’s short-term funds, in a covered interest arbitrage with Japan. Using the following quotes, can Derek make a covered interest arbitrage profit? If so, show the steps and calculate the amount of profit in USD.

Arbitrage funds available

$1,000,000

Spot exchange rate (¥/$)

¥106.00/$

6-month forward rate (¥/$)

¥103.50/$

US dollar 6-month interest rate

4%

Japanese yen 6-month interest rate

2%
Answer #1
This a problem on Interest Rate parity -
It says that - F/S = (1+ia)/(1+ib)
F = Forward Rate
S = Spot Rate
ia = Interest rate of A
ib = Interest rate on B
If this equation holds good then the exchange rate across the globe will be the same.
Yen ia = 2%
Dollor ib = 4%
s = Yen/$ 106
F = Yen/$ 103.5
Calculate the F as per IRP =
F/106 = 1.02^(1/2)/1.04^(1/2) (Assuming compounded annually)
F = 1.0198 x 106
F = 108.0992
Here the Should be Fwd rate is not equal to the actual fwd rate, there is a opportunity for arbitrage.
Stepts in Arbitrage -
Step 1 = Borrow $ 1000000 today for 6 months at 4%
$ payable at the end of 6 months = 1000000 x (1.04^(1/2)) = $        10,19,803.90
Step 2 = Convert $ to Yen using exchange rate 106 today.
1000000 x 106 = ¥ 10,60,00,000.00
Step 3 = Invest these Yen for 6 months at 2% rate.
Yen Inflow on maturity = 106000000 x 1.02^(1/2)
€ 10,70,54,752.35
Step 4 = Cover this amount using forward rate-
Sell Yen 6 months forward at 103.5
$ receivable at the end of 6 months =
107054752.35/103.5 = $        10,34,345.43
Profit = $ Receivable $        10,34,345.43
- $ payable $        10,19,803.90
Profit - $              14,541.53